I need help with the following:
Find p(3a-5)
P(x)=-6x-10
Answer: p(3a-5)=

Answer by jsmallt9(3759)

p(x) = -6x-10
These problems become very easy once you get what the equation above is telling you. In that equation, the "x" in the parentheses is just a placeholder for whatever input you happen to provide to the function p. The rights side of the equation tells us what function p will do with its input (represented by the "x"): It will multiply the input by -6 and then subtract 10. Function p will do this with every input. It will multiply the input by -6 and subtract 10.

p(3a-5) has "3a-5" in the parentheses, the position for the input to the function. p(3a-5) says "Use an input of 3a-5 in function p". And what does function p do with its input? It multiplies the input by -6 and then subtracts 10!
p(3a-5) = -6(3a-5)-10 (Note the parentheses! Very important.)
Now we simplify:
p(3a-5) = -18a+30-10
p(3a-5) = -18a+20